### v1

@@ 3.5 Optimization properties
We conclude here with several properties that can be useful for various extensions and gradient-based approaches. Recall the binwise envelope ${\mathcal{H}}_{b}(q;\bar{p},m)$ and the per-cluster factor $\Phi_{\ell}(q)=\prod_{j=1}^{d}\bigl[{\mathcal{H}}_{b_{\ell j}^{\star}}(q;\bar{p}_{\ell j},m_{\ell})\bigr]^{w_{\ell j}}$, with $w_{\ell j}=m_{\ell j}/m_{\ell}$, $q=s_{i}$, $c_{\ell j}\stackrel{{\scriptstyle\text{def}}}{{=}}\tfrac{q}{b_{\ell j}^{\star}}+1$.
@@
Fix $m\in\mathbb{N}$, $b>0$, $q>0$, $c\stackrel{{\scriptstyle\text{def}}}{{=}}q/b+1>1$, and $z\in[0,1]$.
@@ 4 Related Work
The reader familiar with Spectral Clustering wonders why the graph Laplacian does not appear in our soft assignment formulation. Let ${\bm{f}}\in\{0,1\}^{n}$ denote the indicator of a cluster $\mathbb{C}$}, and let ${\bm{p}}\in[0,1]^{n}$ with $p_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\Pr[{\bm{f}}_{i}=1]$. Then:
@@ 4.2 SimCLR as a special case of our framework
${\bm{W}}_{ab}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\kappa\!\bigl(z_{a},z_{b}\bigr),$
@@ 4.3 CLIP as a probabilistic graph cut
${\bm{W}}_{ij}^{xt}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\kappa\!\bigl(z_{i},u_{j}\bigr),\qquad{\bm{W}}=\begin{pmatrix}0&{\bm{W}}^{xt}\\ ({\bm{W}}^{xt})^{\top}&0\end{pmatrix}$.
@@ A.1 Generalized Poisson-Binomial
Let $r_{i}\sim\mathrm{Bernoulli}(\alpha_{i})$ be independent and define $X\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i=1}^{m}\beta_{i}r_{i}$ with $\beta_{i}\in\mathbb{Z}_{\geq 0}$. Then the probability generating function $G_{X}(t)\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}[t^{X}]$ (for $|t|\leq 1$) is
@@
${\mathcal{I}}(q,{\bm{\alpha}},\beta)\stackrel{{\scriptstyle\text{def}}}{{=}}\int_{0}^{1}\!\Biggl[\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta}\bigr)\Biggr]t^{q-1}\,dt.$
@@
$\sum_{i=1}^{m}\log\!\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta}\bigr)\;\leq\;m\log\!\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr),\quad\bar{\alpha}\stackrel{{\scriptstyle\text{def}}}{{=}}\tfrac{1}{m}\sum_{i=1}^{m}\alpha_{i}.$
@@
Fix $\tau\in[0,1]$ and set $c\stackrel{{\scriptstyle\text{def}}}{{=}}1-\tau$. Define $f_{\tau}(\alpha)\stackrel{{\scriptstyle\text{def}}}{{=}}\log(1-\alpha+\alpha\tau)=\log(1-c\,\alpha)$, a concave function on $[0,1]$ with:
@@
$I(q;{\bm{\alpha}},\beta)\ \stackrel{{\scriptstyle\text{def}}}{{=}}\ \int_{0}^{1}\!\Biggl[\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta}\bigr)\Biggr]t^{q-1}\,dt,\quad B_{\mathrm{AMGM}}(q)\ \stackrel{{\scriptstyle\text{def}}}{{=}}\ \int_{0}^{1}\!\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}t^{q-1}\,dt,$
@@
$\underline{\Delta}(q)\stackrel{{\scriptstyle\text{def}}}{{=}}\int_{0}^{1}\!t^{q-1}\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}\Bigl(1-e^{-\frac{m}{2}\,\gamma(t)\,\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
$\overline{\Delta}(q)\stackrel{{\scriptstyle\text{def}}}{{=}}\int_{0}^{1}\!t^{q-1}\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}\Bigl(1-e^{-\frac{m}{2}\,L(t)\,\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
where $\gamma(t)\stackrel{{\scriptstyle\text{def}}}{{=}}(1-t^{\beta})^{2}$ and $L(t)\stackrel{{\scriptstyle\text{def}}}{{=}}\gamma(t)/t^{2\beta}$, and $B_{\mathrm{AMGM}}(q)\stackrel{{\scriptstyle\text{def}}}{{=}}{\mathcal{H}}_{\beta}(q,\bar{\alpha},\beta)$.
@@ A.5 Proofs for 0-aware gap
Fix $t\in[0,1]$ and write $\tau\stackrel{{\scriptstyle\text{def}}}{{=}}t^{\beta}\in[0,1]$. Let $f_{\tau}(\alpha)\stackrel{{\scriptstyle\text{def}}}{{=}}\log(1-\alpha+\alpha\tau)$, which is concave on $[0,1]$ with
@@
Set the zero-aware weights $\lambda_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\omega_{0}(\alpha_{i})/\Omega$ and denote the $\omega_{0}$–weighted mean and variance by $\bar{\alpha}^{\omega_{0}}=\sum_{i}\lambda_{i}\alpha_{i}$ and $\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})=\sum_{i}\lambda_{i}(\alpha_{i}-\bar{\alpha}^{\omega_{0}})^{2}$ (with the usual convention $\mathrm{Var}^{\omega_{0}}=0$ if $\Omega=0$).
@@
$f_{\tau}(\bar{\alpha}^{\omega_{0}})-\sum_{i=1}^{m}\lambda_{i}f_{\tau}(\alpha_{i})\ \leq\ \frac{L_{\tau}}{2}\,\mathrm{Var}^{\omega_{0}}({\bm{\alpha}}),\qquad L_{\tau}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{(1-\tau)^{2}}{\tau^{2}}.$
@@
$B_{\mathrm{AMGM}}(q)-{\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\ \leq\ \frac{m}{2}\,\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})\;\widetilde{{\mathcal{A}}}\bigl(q;\bar{\alpha},m\bigr)\ \stackrel{{\scriptstyle\text{def}}}{{=}}\ {\mathbb{A}}(q,{\bm{\alpha}},m),$
@@ A.6 Hölder product bound for heterogeneous exponents
Let ${\bm{\beta}}=(\beta_{i})_{i=1}^{m}$ and take $d$ distinct values $\{b_{1},\dots,b_{d}\}\subset(0,\infty)$, and partition indices by $S_{j}\stackrel{{\scriptstyle\text{def}}}{{=}}\{i:\beta_{i}=b_{j}\}$ with sizes $m_{j}\stackrel{{\scriptstyle\text{def}}}{{=}}|S_{j}|$ and $\sum_{j=1}^{d}m_{j}=m$. Define the group means:
$\bar{\alpha}_{j}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{1}{m_{j}}\sum_{i\in S_{j}}\alpha_{i}\in[0,1].$
@@
${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\ \stackrel{{\scriptstyle\text{def}}}{{=}}\ \int_{0}^{1}\Biggl[\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta_{i}}\bigr)\Biggr]\,t^{q-1}\,dt,\qquad q>0.$
@@
$P_{j}(t)\ \stackrel{{\scriptstyle\text{def}}}{{=}}\ \prod_{i\in S_{j}}\bigl(1-\alpha_{i}+\alpha_{i}t^{b_{j}}\bigr)\ \leq\ \bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m_{j}}\quad\text{by AM--GM.}$
@@
Let $w_{j}\stackrel{{\scriptstyle\text{def}}}{{=}}m_{j}/m$ and split $t^{q-1}=\prod_{j=1}^{d}t^{(q-1)w_{j}}$. Set:
$g_{j}(t)\ \stackrel{{\scriptstyle\text{def}}}{{=}}\ \bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m_{j}}\,t^{(q-1)w_{j}}.$
Choose exponents $p_{j}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{m}{m_{j}}>1$, so that $\sum_{j=1}^{d}\frac{1}{p_{j}}=\sum_{j}\frac{m_{j}}{m}=1$. By Hölder’s inequality for products,
@@ B.4 Gradient of the Hölder envelope for heterogeneous $\beta$
Recall the Hölder envelope (Sec. A.6): with distinct exponents $\{b_{1},\dots,b_{d}\}$, groups $S_{k}\!\stackrel{{\scriptstyle\text{def}}}{{=}}\!\{i:\beta_{i}=b_{k}\}$, sizes $m_{k}\!=\!|S_{k}|$, $m=\sum_{k}m_{k}$, and means $\bar{\alpha}_{k}\!=\!\frac{1}{m_{k}}\sum_{i\in S_{k}}\alpha_{i}$, we defined:
$B_{\mathrm{Holder}}\;=\;\prod_{k=1}^{d}\bigl(B_{k}^{\ast}\bigr)^{m_{k}/m},\qquad B_{k}^{\ast}\;=\;\frac{1}{q}\;{}_{2}F_{1}\!\Bigl(-m,\,1;\,c_{k};\,\bar{\alpha}_{k}\Bigr),\quad c_{k}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{q}{b_{k}}+1,$
@@
Let $F_{1}(z)\!\stackrel{{\scriptstyle\text{def}}}{{=}}\!{}_{2}F_{1}(-m,1;c_{j};z)$ and $F_{2}(z)\!\stackrel{{\scriptstyle\text{def}}}{{=}}\!{}_{2}F_{1}(-m+1,2;c_{j}+1;z)$. By the derivative identity, $F_{2}(z)=-\frac{c_{j}}{m}\,\frac{d}{dz}F_{1}(z)$, hence
@@ B.5 Gradients of the final objective
For cluster $\ell$ and bin index $j$, let $S_{\ell j}$ be the set of vertices assigned to bin $j$ (with common exponent $b_{j}$), $m_{\ell j}\stackrel{{\scriptstyle\text{def}}}{{=}}|S_{\ell j}|$, $m_{\ell}\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{j}m_{\ell j}$, and:
$\bar{p}_{\ell j}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{1}{m_{\ell j}}\sum_{r\in S_{\ell j}}p_{r\ell},\qquad w_{\ell j}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{m_{\ell j}}{m_{\ell}}\quad(\text{Hölder weight}).$
@@
${\mathcal{A}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell})\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\sum_{r=0}^{2}\binom{2}{r}(-1)^{r}\,{\mathcal{H}}_{b_{j}}\!\bigl(q+rb_{j};\bar{p}_{\ell j},m_{\ell}\bigr),\qquad\widetilde{{\mathcal{A}}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell})\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{\partial}{\partial\bar{p}_{\ell j}}{\mathcal{A}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell}).$
@@
$\omega_{i}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\omega(p_{i\ell}),\qquad\Omega_{\ell j}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{r\in S_{\ell j}}\omega(p_{r\ell}),\qquad S_{2}\;\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{r\in S_{\ell j}}\omega(p_{r\ell})\,p_{r\ell},$
$\mu\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\bar{p}^{\omega}_{\ell j}\;=\;S_{2}/\Omega_{\ell j},\qquad V\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\mathrm{Var}^{\omega}_{\ell j}(p)\;=\;\frac{1}{\Omega_{\ell j}}\sum_{r\in S_{\ell j}}\omega(p_{r\ell})\bigl(p_{r\ell}-\mu\bigr)^{2},$
with the convention $V=0$ if $\Omega_{\ell j}=0$. In the paper we take $\omega(x)=x(1-x)$ so that $\omega^{\prime}_{i}\stackrel{{\scriptstyle\text{def}}}{{=}}\frac{d}{dp}\omega(p_{i\ell})=1-2p_{i\ell}$ (zero-aware and symmetric). The derivatives of the weighted mean and variance are:
@@
Let $q_{r}\stackrel{{\scriptstyle\text{def}}}{{=}}q+rb_{j}$ and $c_{r}\stackrel{{\scriptstyle\text{def}}}{{=}}\tfrac{q_{r}}{b_{j}}+1$. Using $\tfrac{d}{dz}{}_{2}F_{1}(a,b;c;z)=\frac{ab}{c}{}_{2}F_{1}(a+1,b+1;c+1;z)$,
@@
$\Gamma^{\mathrm{ewa}}_{\ell j}(q)\;\stackrel{{\scriptstyle\text{def}}}{{=}}\;\frac{m_{\ell}}{2}\,w_{\ell j}\;V\;\widetilde{{\mathcal{A}}}_{b_{j}}\!\bigl(q;\bar{p}_{\ell j},m_{\ell}\bigr),$

### v2

@@ 3.5 Optimization properties
We conclude here with several properties that can be useful for various extensions and gradient-based approaches. Recall the binwise envelope ${\mathcal{H}}_{b}(q;\bar{p},m)$ and the per-cluster factor $\Phi_{\ell}(q)=\prod_{j=1}^{d}\bigl[{\mathcal{H}}_{b_{\ell j}^{\star}}(q;\bar{p}_{\ell j},m_{\ell})\bigr]^{w_{\ell j}}$, with $w_{\ell j}=m_{\ell j}/m_{\ell}$, $q=s_{i}$, $c_{\ell j}\vcentcolon=\tfrac{q}{b_{\ell j}^{\star}}+1$.
@@
Fix $m\in\mathbb{N}$, $b>0$, $q>0$, $c\vcentcolon=q/b+1>1$, and $z\in[0,1]$.
@@ 4 Related Work
The reader familiar with Spectral Clustering wonders why the graph Laplacian does not appear in our soft assignment formulation. Let ${\bm{f}}\in\{0,1\}^{n}$ denote the indicator of a cluster $\mathbb{C}$}, and let ${\bm{p}}\in[0,1]^{n}$ with $p_{i}\vcentcolon=\Pr[{\bm{f}}_{i}=1]$. Then:
@@ 4.2 SimCLR as a special case of our framework
${\bm{W}}_{ab}\;\vcentcolon=\;\kappa\!\bigl(z_{a},z_{b}\bigr),$
@@ 4.3 CLIP as a probabilistic graph cut
${\bm{W}}_{ij}^{xt}\;\vcentcolon=\;\kappa\!\bigl(z_{i},u_{j}\bigr),\qquad{\bm{W}}=\begin{pmatrix}0&{\bm{W}}^{xt}\\ ({\bm{W}}^{xt})^{\top}&0\end{pmatrix}$.
@@ A.1 Generalized Poisson-Binomial
Let $r_{i}\sim\mathrm{Bernoulli}(\alpha_{i})$ be independent and define $X\vcentcolon=\sum_{i=1}^{m}\beta_{i}r_{i}$ with $\beta_{i}\in\mathbb{Z}_{\geq 0}$. Then the probability generating function $G_{X}(t)\vcentcolon=\mathbb{E}[t^{X}]$ (for $|t|\leq 1$) is
@@
${\mathcal{I}}(q,{\bm{\alpha}},\beta)\vcentcolon=\int_{0}^{1}\!\Biggl[\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta}\bigr)\Biggr]t^{q-1}\,dt.$
@@
$\sum_{i=1}^{m}\log\!\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta}\bigr)\;\leq\;m\log\!\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr),\quad\bar{\alpha}\vcentcolon=\tfrac{1}{m}\sum_{i=1}^{m}\alpha_{i}.$
@@
Fix $\tau\in[0,1]$ and set $c\vcentcolon=1-\tau$. Define $f_{\tau}(\alpha)\vcentcolon=\log(1-\alpha+\alpha\tau)=\log(1-c\,\alpha)$, a concave function on $[0,1]$ with:
@@
$I(q;{\bm{\alpha}},\beta)\ \vcentcolon=\ \int_{0}^{1}\!\Biggl[\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta}\bigr)\Biggr]t^{q-1}\,dt,\quad B_{\mathrm{AMGM}}(q)\ \vcentcolon=\ \int_{0}^{1}\!\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}t^{q-1}\,dt,$
@@
$\underline{\Delta}(q)\vcentcolon=\int_{0}^{1}\!t^{q-1}\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}\Bigl(1-e^{-\frac{m}{2}\,\gamma(t)\,\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
$\overline{\Delta}(q)\vcentcolon=\int_{0}^{1}\!t^{q-1}\bigl(1-\bar{\alpha}+\bar{\alpha}\,t^{\beta}\bigr)^{m}\Bigl(1-e^{-\frac{m}{2}\,L(t)\,\mathrm{Var}({\bm{\alpha}})}\Bigr)\,dt,$
where $\gamma(t)\vcentcolon=(1-t^{\beta})^{2}$ and $L(t)\vcentcolon=\gamma(t)/t^{2\beta}$, and $B_{\mathrm{AMGM}}(q)\vcentcolon={\mathcal{H}}_{\beta}(q,\bar{\alpha},\beta)$.
@@ A.5 Proofs for 0-aware gap
Fix $t\in[0,1]$ and write $\tau\vcentcolon=t^{\beta}\in[0,1]$. Let $f_{\tau}(\alpha)\vcentcolon=\log(1-\alpha+\alpha\tau)$, which is concave on $[0,1]$ with
@@
Set the zero-aware weights $\lambda_{i}\vcentcolon=\omega_{0}(\alpha_{i})/\Omega$ and denote the $\omega_{0}$–weighted mean and variance by $\bar{\alpha}^{\omega_{0}}=\sum_{i}\lambda_{i}\alpha_{i}$ and $\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})=\sum_{i}\lambda_{i}(\alpha_{i}-\bar{\alpha}^{\omega_{0}})^{2}$ (with the usual convention $\mathrm{Var}^{\omega_{0}}=0$ if $\Omega=0$).
@@
$f_{\tau}(\bar{\alpha}^{\omega_{0}})-\sum_{i=1}^{m}\lambda_{i}f_{\tau}(\alpha_{i})\ \leq\ \frac{L_{\tau}}{2}\,\mathrm{Var}^{\omega_{0}}({\bm{\alpha}}),\qquad L_{\tau}\vcentcolon=\frac{(1-\tau)^{2}}{\tau^{2}}.$
@@
$B_{\mathrm{AMGM}}(q)-{\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\ \leq\ \frac{m}{2}\,\mathrm{Var}^{\omega_{0}}({\bm{\alpha}})\;\widetilde{{\mathcal{A}}}\bigl(q;\bar{\alpha},m\bigr)\ \vcentcolon=\ {\mathbb{A}}(q,{\bm{\alpha}},m),$
@@ A.6 Hölder product bound for heterogeneous exponents
Let ${\bm{\beta}}=(\beta_{i})_{i=1}^{m}$ and take $d$ distinct values $\{b_{1},\dots,b_{d}\}\subset(0,\infty)$, and partition indices by $S_{j}\vcentcolon=\{i:\beta_{i}=b_{j}\}$ with sizes $m_{j}\vcentcolon=|S_{j}|$ and $\sum_{j=1}^{d}m_{j}=m$. Define the group means:
$\bar{\alpha}_{j}\;\vcentcolon=\;\frac{1}{m_{j}}\sum_{i\in S_{j}}\alpha_{i}\in[0,1].$
@@
${\mathcal{I}}(q,{\bm{\alpha}},{\bm{\beta}})\ \vcentcolon=\ \int_{0}^{1}\Biggl[\prod_{i=1}^{m}\bigl(1-\alpha_{i}+\alpha_{i}t^{\beta_{i}}\bigr)\Biggr]\,t^{q-1}\,dt,\qquad q>0.$
@@
$P_{j}(t)\ \vcentcolon=\ \prod_{i\in S_{j}}\bigl(1-\alpha_{i}+\alpha_{i}t^{b_{j}}\bigr)\ \leq\ \bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m_{j}}\quad\text{by AM--GM.}$
@@
Let $w_{j}\vcentcolon=m_{j}/m$ and split $t^{q-1}=\prod_{j=1}^{d}t^{(q-1)w_{j}}$. Set:
$g_{j}(t)\ \vcentcolon=\ \bigl(1-\bar{\alpha}_{j}+\bar{\alpha}_{j}t^{b_{j}}\bigr)^{m_{j}}\,t^{(q-1)w_{j}}.$
Choose exponents $p_{j}\vcentcolon=\frac{m}{m_{j}}>1$, so that $\sum_{j=1}^{d}\frac{1}{p_{j}}=\sum_{j}\frac{m_{j}}{m}=1$. By Hölder’s inequality for products,
@@ B.4 Gradient of the Hölder envelope for heterogeneous $\beta$
Recall the Hölder envelope (Sec. A.6): with distinct exponents $\{b_{1},\dots,b_{d}\}$, groups $S_{k}\!\vcentcolon=\!\{i:\beta_{i}=b_{k}\}$, sizes $m_{k}\!=\!|S_{k}|$, $m=\sum_{k}m_{k}$, and means $\bar{\alpha}_{k}\!=\!\frac{1}{m_{k}}\sum_{i\in S_{k}}\alpha_{i}$, we defined:
$B_{\mathrm{Holder}}\;=\;\prod_{k=1}^{d}\bigl(B_{k}^{\ast}\bigr)^{m_{k}/m},\qquad B_{k}^{\ast}\;=\;\frac{1}{q}\;{}_{2}F_{1}\!\Bigl(-m,\,1;\,c_{k};\,\bar{\alpha}_{k}\Bigr),\quad c_{k}\vcentcolon=\frac{q}{b_{k}}+1,$
@@
Let $F_{1}(z)\!\vcentcolon=\!{}_{2}F_{1}(-m,1;c_{j};z)$ and $F_{2}(z)\!\vcentcolon=\!{}_{2}F_{1}(-m+1,2;c_{j}+1;z)$. By the derivative identity, $F_{2}(z)=-\frac{c_{j}}{m}\,\frac{d}{dz}F_{1}(z)$, hence
@@ B.5 Gradients of the final objective
For cluster $\ell$ and bin index $j$, let $S_{\ell j}$ be the set of vertices assigned to bin $j$ (with common exponent $b_{j}$), $m_{\ell j}\vcentcolon=|S_{\ell j}|$, $m_{\ell}\vcentcolon=\sum_{j}m_{\ell j}$, and:
$\bar{p}_{\ell j}\;\vcentcolon=\;\frac{1}{m_{\ell j}}\sum_{r\in S_{\ell j}}p_{r\ell},\qquad w_{\ell j}\;\vcentcolon=\;\frac{m_{\ell j}}{m_{\ell}}\quad(\text{Hölder weight}).$
@@
${\mathcal{A}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell})\;\vcentcolon=\;\sum_{r=0}^{2}\binom{2}{r}(-1)^{r}\,{\mathcal{H}}_{b_{j}}\!\bigl(q+rb_{j};\bar{p}_{\ell j},m_{\ell}\bigr),\qquad\widetilde{{\mathcal{A}}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell})\;\vcentcolon=\;\frac{\partial}{\partial\bar{p}_{\ell j}}{\mathcal{A}}_{b_{j}}(q;\bar{p}_{\ell j},m_{\ell}).$
@@
$\omega_{i}\;\vcentcolon=\;\omega(p_{i\ell}),\qquad\Omega_{\ell j}\;\vcentcolon=\sum_{r\in S_{\ell j}}\omega(p_{r\ell}),\qquad S_{2}\;\vcentcolon=\sum_{r\in S_{\ell j}}\omega(p_{r\ell})\,p_{r\ell},$
$\mu\;\vcentcolon=\;\bar{p}^{\omega}_{\ell j}\;=\;S_{2}/\Omega_{\ell j},\qquad V\;\vcentcolon=\;\mathrm{Var}^{\omega}_{\ell j}(p)\;=\;\frac{1}{\Omega_{\ell j}}\sum_{r\in S_{\ell j}}\omega(p_{r\ell})\bigl(p_{r\ell}-\mu\bigr)^{2},$
with the convention $V=0$ if $\Omega_{\ell j}=0$. In the paper we take $\omega(x)=x(1-x)$ so that $\omega^{\prime}_{i}\vcentcolon=\frac{d}{dp}\omega(p_{i\ell})=1-2p_{i\ell}$ (zero-aware and symmetric). The derivatives of the weighted mean and variance are:
@@
Let $q_{r}\vcentcolon=q+rb_{j}$ and $c_{r}\vcentcolon=\tfrac{q_{r}}{b_{j}}+1$. Using $\tfrac{d}{dz}{}_{2}F_{1}(a,b;c;z)=\frac{ab}{c}{}_{2}F_{1}(a+1,b+1;c+1;z)$,
@@
$\Gamma^{\mathrm{ewa}}_{\ell j}(q)\;\vcentcolon=\;\frac{m_{\ell}}{2}\,w_{\ell j}\;V\;\widetilde{{\mathcal{A}}}_{b_{j}}\!\bigl(q;\bar{p}_{\ell j},m_{\ell}\bigr),$